OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? x -2 0 4 f(x) -7 3 -73 A. y = -3x^2 - 3x + 4 B. y = 4x^2 + 3x - 3 C. y = -4x^2 - 3x + 3 D. y = -3x^2 - 4x + 3 my answer is B can some one tell me if im correct on this

jimthompson5910 (jim_thompson5910):

it can't be B, here's why when x = -2, y is.. y = 4x^2 + 3x - 3 y = 4(-2)^2 + 3(-2) - 3 y = 4(4) + 3(-2) - 3 y = 16 - 6 - 3 y = 7 which is NOT -7 as shown in the table above.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be C

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

good i thought that was it but for some reason i kept thinking it was B